OpenStudy (anonymous):

A(f) = 1 + f/30 )what is the inverse?

OpenStudy (anonymous):

is the answer; a/30 - 1 ?

OpenStudy (anonymous):

A^-1 (f) = 30 (f - 1)

OpenStudy (anonymous):

im so confused :(

OpenStudy (anonymous):

Write it as A(f) = (30 + f)/30 and then go from there.

OpenStudy (anonymous):

Did you understand it yet ?

OpenStudy (anonymous):

Or you want me to elaborate it ?

OpenStudy (anonymous):

elaborate please

OpenStudy (anonymous):

ok give me a sec.

OpenStudy (anonymous):

Replace A(f) with "y" so it will be y = 1 + f/30 Now switch "y" with "f" f = 1 + y/30 multiply by 30 on both sides 30f = 30 + y Now 30f -30 = 30 +y -30 which leaves us 30f -30 = y Now y = 30 (f - 1) Replace "y" with A^-1(f)

OpenStudy (anonymous):

ok I believe I got it ... thanks

OpenStudy (anonymous):

Welcome.

OpenStudy (anonymous):

wait sorry ....I don't understand the - 30 in front of y

OpenStudy (anonymous):

-30 on both sides to remove the +30 with "y", Because you can not just do -30 on one side. The equation has to be balanced.

OpenStudy (anonymous):

Or you can just do this 30f = 30 + y 30f - 30 = y and then continue.

OpenStudy (anonymous):

Thank you !!